(x + 7)(x - 2) = -20
x =

2b(4b + 5) = 0
b =

enter your answers as a comma-separated list
Answer #1

(x + 7)(x - 2) = -20

x2-2x+7x-14 = -20

x2+5x+6 = 0

x2+2x+3x+6 = 0

x(x+2)+3(x+2) = 0

(x+2)(x+3) = 0

x=-2, x= -3

x = 2b(4b + 5)

if  x=-2

8b2+10b+2 = 0

8b2+8b+2b+2 =0

8b(b+1)+2(b+1)=0

(b+1)(8b+2)=0

b=-1/4, b=-1

if  x=-3

8b2+10b+3 = 0

8b2+4b+6b+3 =0

4b(2b+1)+3(2b+1)=0

(2b+1)(4b+3) = 0

b = -1/2 , b = -3/4

So value of b = -1,-1/4,-1/2,-3/4.

Please provide feedback.Thumbs up.Thanks.